OpenStudy (anonymous):

solve the following systems of equations algebraically 55: y=-7x+15 y=-5x+11

OpenStudy (anonymous):

Algebraically is a pretty vague adjective, in the context of solving a system of linear of equations.

OpenStudy (anonymous):

You have two equations equalling y. Therefore both RHS of the two equations must be equal. -7x+15=-5x+11 4=2x 2=x Not done yet...still need y. We have x=2, now let's plug this in to either equation (we will get same answer either way) y = -7x+15 = -7(2)+15 = -14 + 15 y = 1 So x = 2, y = 1

OpenStudy (anonymous):

thanks :)